### v1

# Graph-Structured Driven Dual Adaptation for Mitigating Popularity Bias

Abstract
Popularity bias is a common challenge in recommender systems. It often causes unbalanced item recommendation performance and intensifies the Matthew effect. Due to limited user-item interactions, unpopular items are frequently constrained to the embedding neighborhoods of only a few users, leading to representation collapse and weakening the model’s generalization. Although existing supervised alignment and reweighting methods can help mitigate this problem, they still face two major limitations: (1) they overlook the inherent variability among different Graph Convolutional Networks(GCNs) layers, which can result in negative gains in deeper layers; (2) they rely heavily on fixed hyperparameters to balance popular and unpopular items, limiting adaptability to diverse data distributions and increasing model complexity.
To address these issues, we propose  Graph-Structured Dual Adaptation Framework (GSDA). First, our theoretical analysis reveals a critical limitation of supervised alignment strategies arising from the over-smoothing effect in GCNs. Specifically, as GCNs layers deepen, the distinctiveness between popular and unpopular items gradually diminishes, which can be quantitatively measured by the reduction in their conditional entropy. This progressive loss of distinctiveness severely weakens the effectiveness of supervised alignment methods in alleviating popularity bias at deeper layers. Motivated by this theoretical insight,  GSDA  simultaneously mines structural characteristics and distribution information from the graph adjacency matrix and incorporates a dual adaptive strategy. First, the hierarchical adaptive alignment mechanism uses the F-norm of the adjacency matrix to perform layer-specific weight decay, mitigating the negative effects of deep conditional entropy decay. Second, the distribution-aware dynamic contrast weighting strategy, based on a real-time Gini coefficient, removes the dependence on fixed hyperparameters and enables the model to adapt to diverse data distributions. Extensive experiments on three benchmark datasets show that  GSDA  not only significantly alleviates popularity bias but also consistently outperforms state-of-the-art methods in recommendation performance.
Index Terms:
Recommender Systems, Popularity Bias, Supervised Alignment, Re-weighting, Contrastive Learning
I Introduction
Figure 1: Illustration of representation collapse. The popular item ($i_{1}$) interacts with multiple users ($u_{1},u_{2},u_{3}$), obtaining a generalized representation, while the unpopular item ($i_{2}$), interacting with only one user ($u_{2}$), suffers from representation collapse due to insufficient contextual information.
Recommender systems (RS) have become indispensable in various domains, such as e-commerce and streaming services, providing personalized recommendations to alleviate information overload [1, 2, 3, 4, 5]. Among numerous recommendation paradigms, Collaborative Filtering (CF) remains a cornerstone approach due to its ability to achieve strong recommendation performance by solely leveraging historical user-item interactions. Recently, Graph Convolutional Networks (GCNs) have emerged as prominent methods in CF, effectively capturing high-order collaborative signals through graph structures. However, popularity bias persists as a critical challenge, wherein highly-interacted items dominate recommendation lists, causing unpopular items to be increasingly marginalized [6, 7, 8]. Such imbalance not only skews recommendation outcomes but also intensifies the ”Matthew effect” [9, 10, 11, 12], further widening the visibility gap between popular and unpopular items.
The mechanism of popularity bias is closely tied to the alignment paradigm in Collaborative Filtering (CF), where user and item representations are aligned according to observed user-item interactions [13, 14, 15, 16]. However, when these interactions are highly imbalanced, this alignment strategy often results in representation collapse for unpopular items. As illustrated in Fig.1, a target user $u_{4}$ interacts with both a popular item ($i_{1}$) and an unpopular item ($i_{2}$) during the testing phase. During training, the popular item $i_{1}$ connects to multiple users ($u_{1},u_{2},$ and $u_{3}$), enabling it to aggregate richer information and learn a more generalized embedding. In contrast, the unpopular item $i_{2}$ has minimal exposure, relying excessively on the embedding space of a single user ($u_{2}$). Consequently, $i_{2}$ experiences representation collapse due to limited contextual information, further marginalizing unpopular items and exacerbating popularity bias within CF-based recommender systems. This representation collapse fundamentally arises from CF’s supervised alignment paradigm, which relies on user-item interactions for embedding optimization.
To address representation collapse, existing methods mainly focus on improving embedding distributions (e.g., AURL [7]), enhancing embedding uniformity through contrastive learning (e.g., SimGCL [17], SGL [18]), and employing group-specific weighting or supervised alignment to explicitly differentiate item popularity levels [19, 20, 21, 22]. More recently, supervised alignment methods have emerged, explicitly aligning embeddings of popular and unpopular items interacting with the same user [19]. Although these approaches have made considerable progress, they still suffer from critical limitations.
First, existing supervised alignment methods typically overlook the inherent structural differences among layers in Graph Convolutional Networks (GCNs). Specifically, shallow layers focus primarily on capturing local neighborhood features, whereas deeper layers aggregate broader, global structural information. Uniformly applying the same alignment constraints across all layers neglects this hierarchy, resulting in ineffective or even counterproductive alignment in deeper layers due to embedding homogenization and semantic drift. Second, current group-specific weighting strategies heavily rely on static hyperparameters to balance the contributions of popular and unpopular items. Such rigid parameter settings require extensive manual tuning for different datasets, limiting model adaptability and scalability in dynamic environments with varying item popularity distributions.
To overcome these limitations, we propose a novel framework called  Graph-Structured Dual Adaptation Framework (GSDA), which explicitly leverages structural and distributional insights derived from the graph’s adjacency matrix to effectively mitigate popularity bias. Our theoretical analysis reveals a critical yet previously overlooked issue: as Graph Convolutional Networks (GCNs) deepen, the over-smoothing effect leads to the homogenization of node embeddings, which significantly reduces the conditional entropy between popular and unpopular items. This reduction in conditional entropy severely compromises the effectiveness of supervised alignment strategies in deeper layers. Motivated by this novel theoretical insight, we propose a dual adaptive strategy that uniquely combines layer-wise structural signals and item distribution information from the adjacency matrix. Specifically, our approach  GSDA  includes: (1) a hierarchical adaptive alignment mechanism to address embedding collapse across layers; (2) a distribution-aware dynamic weighting strategy that avoids reliance on fixed hyperparameters by dynamically adapting to changes in item popularity distributions. Extensive experiments validate that our approach effectively mitigates popularity bias and consistently surpasses state-of-the-art baselines in recommendation accuracy. Our main contributions are summarized as follows:
  • We theoretically and empirically reveal that supervised alignment effectiveness declines significantly in deeper GCN layers due to embedding homogenization (over-smoothing) induced by repeated propagation through higher-order adjacency matrices.
  • We introduce a dual adaptive framework that leverages graph structural information to extract item distribution patterns and dynamically adjusts alignment strategies, effectively mitigating popularity bias without extensive hyperparameter tuning.
  • Extensive experiments demonstrate our approach’s superior performance, significantly reducing popularity bias while achieving state-of-the-art recommendation accuracy and robustness across diverse scenarios.
II Preliminary
II-A Recommendation Task
Recommender systems aim to predict user preferences toward unobserved items based on historical user-item interactions. Collaborative Filtering (CF), a widely adopted paradigm, predicts these interactions by aligning user and item representations. User-item interactions are represented by a binary matrix $\mathcal{R}\in\{0,1\}^{M\times N}$, where $\mathcal{R}_{u,i}=1$ indicates an interaction between user $u$ and item $i$, and 00 otherwise. Let $U$ and $I$ denote user and item sets with $|U|=M$ and $|I|=N$, respectively. CF models typically learn user embeddings $\mathbf{Z}\in\mathbb{R}^{M\times D}$ and item embeddings $\mathbf{E}\in\mathbb{R}^{N\times D}$, where $D$ is the embedding dimension. The predicted preference score for user $u$ towards item $i$ is calculated as $s(u,i)=\mathbf{z}_{u}^{\top}\mathbf{e}_{i}$, where $\mathbf{z}_{u}$ and $\mathbf{e}_{i}$ are the embeddings for user $u$ and item $i$.
To optimize user and item embeddings, we employ the Bayesian Personalized Ranking (BPR) loss [14], a pairwise ranking objective widely used in recommendation tasks. This loss encourages the model to rank observed (positive) interactions higher than unobserved (negative) interactions:
$\mathcal{L}_{rec}=-\frac{1}{|\mathcal{R}|}\sum_{(u,i,j)\in\mathcal{O^{+}}}ln\sigma(s(u,i)-s(u,j)),$ (1)
where $\sigma(\cdot)$ is the sigmoid function, $\mathcal{O^{+}}=\{(u,i,j)|\mathcal{R}_{u,i}=1,\mathcal{R}_{u,j}=0\}$ represents pairwise data, and $j$ is a randomly sampled negative item that the user has not interacted with.
II-B Graph Convolutional Networks
Graph Convolutional Networks (GCNs) have been widely adopted in collaborative filtering due to their capability of modeling high-order collaborative signals between users and items through graph-based information propagation [23, 24, 25]. Specifically, user-item interactions can be naturally represented as a bipartite graph $\mathcal{G}=(\mathcal{U}\cup\mathcal{I},\mathbf{A})$, where nodes include users and items, and edges represent interactions. The corresponding adjacency matrix $\mathbf{A}\in\mathbb{R}^{(M+N)\times(M+N)}$ is defined as:
$\mathbf{A}=\left[\begin{array}[]{cc}\mathbf{0}^{M\times M}&\mathcal{R}\\[6.0pt] \mathcal{R}^{\top}&\mathbf{0}^{N\times N}\end{array}\right],$ (4)
where $\mathcal{R}\in\{0,1\}^{M\times N}$ denotes the interaction matrix between users and items. To explicitly preserve self-information during the propagation, self-loops are introduced by integrating an identity matrix $\mathbf{I}$, resulting in the adjacency matrix $(\mathbf{A}+\mathbf{I})$. This adjacency matrix is then normalized as follows:
$\hat{\mathbf{A}}=\mathbf{D}^{-1/2}\,(\mathbf{A}+\mathbf{I})\,\mathbf{D}^{-1/2},$ (5)
where $\mathbf{D}$ is the diagonal degree matrix. And its diagonal elements $\mathbf{D}_{ii}$ represent the degree (sum of connections) of node $i$. This normalized adjacency matrix serves as the basis for propagating embeddings within GCN layers [26].
In this work, we employ LightGCN [23] as the embedding encoder due to its simplicity and strong performance in collaborative filtering tasks. Specifically, we initialize embeddings for all nodes (users and items) as $\mathbf{X}^{(0)}=\bigl{[}\mathbf{Z}^{(0)},\mathbf{E}^{(0)}\bigr{]}$, where $\mathbf{Z}^{(0)}$ and $\mathbf{E}^{(0)}$ denote initial embeddings of users and items, respectively. These embeddings are iteratively updated through a layer-wise propagation rule defined as follows:
$\mathbf{X}^{(l)}=\hat{\mathbf{A}}\mathbf{X}^{(l-1)},$ (6)
where $\hat{\mathbf{A}}$ denotes the normalized adjacency matrix, and $\mathbf{X}^{(l-1)}$ represents node embeddings from the previous layer. Applying this propagation recursively across $l$ layers, the final embeddings at layer $l$ can be succinctly expressed as:
$\mathbf{X}^{(l)}=\hat{\mathbf{A}}^{l}\mathbf{X}^{(0)}.$ (7)
It indicates that the embeddings at layer $l$ are determined by the $l$-th order multiplication of the normalized adjacency matrix, thus capturing multi-hop neighborhood information.
III Theoretical and Empirical Validation
In this subsection, we establish a theoretical connection between supervised alignment effectiveness and conditional entropy. Specifically, we first define conditional entropy as a metric for measuring alignment efficacy (Entropy as a Measure of Alignment Effectiveness). Then, we demonstrate how repeated message passing in GCNs leads to over-smoothing, homogenizing embeddings across layers (Over-Smoothing Caused by High-Order Adjacency Matrices). Finally, we analytically derive the layer-wise reduction in conditional entropy due to over-smoothing, showing that it weakens the information transfer capability of supervised alignment at deeper layers (Layer-Wise Conditional Entropy Reduction). This structured analysis provides a comprehensive understanding of why alignment effectiveness diminishes as GCN depth increases.
III-A Theoretical Analyses
Entropy as a Measure of Alignment Effectiveness. Supervised alignment aims to mitigate popularity bias by explicitly aligning embeddings of popular and unpopular items that share interactions with the same user [19]. To quantitatively evaluate the effectiveness of supervised alignment, we leverage an information-theoretic [27, 28] perspective by defining a metric based on entropy. Specifically, we measure how well embeddings of unpopular items retain independent information when conditioned on popular item embeddings. This motivates the use of conditional entropy as a direct indicator of supervised alignment performance.
$H(\mathbf{X}_{up}|\mathbf{X}_{p})=-\mathbb{E}_{P(\mathbf{X}_{up},\mathbf{X}_{p})}[\log P(\mathbf{X}_{up}|\mathbf{X}_{p})],$ (8)
where $\mathbf{X}_{up}$ and $\mathbf{X}_{p}$ represent embeddings of unpopular and popular items, respectively. This entropy-based metric directly reflects alignment efficacy—higher values indicate more effective information transfer. In particular, a higher conditional entropy indicates greater uncertainty about unpopular item embeddings, thus implying more potential for beneficial information transfer from popular to unpopular items. Conversely, lower conditional entropy implies that embeddings of unpopular items already possess similar information to popular items, limiting alignment benefits.
Over-Smoothing Caused by High-Order Adjacency Matrices. Having established entropy as a key metric for alignment effectiveness, we now analyze how GCN message passing affects entropy. In particular, we show that repeated propagation through the normalized adjacency matrix leads to over-smoothing, which directly impacts entropy-based alignment performance. In Graph Convolutional Networks (GCNs), node embeddings at the $l$-th layer are generated by recursively propagating information through the normalized adjacency matrix $\hat{\mathbf{A}}$. Formally, this process is defined as Eq. (7), where $\mathbf{X}^{(0)}$ represents initial node embeddings, typically initialized from a random Gaussian distribution with zero mean, and $\hat{\mathbf{A}}$ is the normalized adjacency matrix encoding structural relationships among nodes.
As the depth $l$ increases, the repeated multiplication of $\hat{\mathbf{A}}$ leads its entries to become progressively uniform [29]. Specifically, the differences between any two entries in the high-order adjacency matrix diminish, satisfying:
$\lim_{l\to\infty}\left|\hat{\mathbf{A}}^{l}_{u,j}-\hat{\mathbf{A}}^{l}_{p,k}\right|=0,\quad\text{for all nodes }u,p,j,k.$ (9)
This convergence indicates that propagation weights between any node pairs become highly similar as layers deepen. Combining this nearly uniform propagation structure with initially random Gaussian embeddings $\mathbf{X}^{(0)}$ results in the homogenization of node embeddings. Such embedding homogenization, known as over-smoothing, significantly reduces the discriminative power of node representations, especially impacting unpopular items with sparse interactions.
Layer-Wise Conditional Entropy Reduction. Given the homogenization of embeddings induced by over-smoothing, we now analyze how this phenomenon directly leads to a monotonic reduction in conditional entropy. By explicitly deriving the relationship between adjacency matrix differences and entropy, we demonstrate that alignment effectiveness inevitably weakens as the number of layers increases. We now explicitly connect the over-smoothing phenomenon described above to the progressive reduction in conditional entropy observed in supervised alignment. Consider the embeddings of an unpopular item $\mathbf{x}_{up}^{(l)}$ and a popular item $\mathbf{x}_{p}^{(l)}$ at layer $l$, which can be expressed based on GCN propagation:
$\mathbf{x}_{up}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{up,j}\mathbf{x}_{j}^{(0)},\quad\mathbf{x}_{p}^{(l)}=\sum_{j}\hat{\mathbf{A}}^{l}_{p,j}\mathbf{x}_{j}^{(0)},$ (10)
where $\hat{\mathbf{A}}^{l}_{up,j}$ and $\hat{\mathbf{A}}^{l}_{p,j}$ represent entries in the normalized adjacency matrix at layer $l$, quantifying how information propagates from node $j$ to the unpopular item $up$ and popular item $p$, respectively. Here, $\mathbf{x}_{j}^{(0)}$ denotes the initial embedding vector of node $j$, initialized from a random Gaussian distribution.
To quantify how well popular item embeddings can inform unpopular item embeddings, we use conditional entropy as defined Eq. (8). For tractable analysis, we approximate the conditional probability $P(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})$ via Gaussian kernel density estimation (KDE) [30]:
$P(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})\;\propto\;\exp\Bigl{(}-\frac{\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}}{2\sigma^{2}}\Bigr{)},$ (11)
where $\sigma$ is a bandwidth parameter controlling the smoothness of the probability estimation. By substituting Eq. (11) into the conditional entropy definition, we explicitly establish a relationship between entropy and embedding distance:
$H(\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)})\;\propto\;\mathbb{E}\bigl{[}\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}\bigr{]}.$ (12)
Expanding the squared Euclidean distance term explicitly using the definitions in Eq. (10) yields:
$\|\mathbf{x}_{up}^{(l)}-\mathbf{x}_{p}^{(l)}\|^{2}\;=\;\sum_{j,k}\,\bigl{(}\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr{)}^{2}\,\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}.$ (13)
This step explicitly clarifies how the differences between adjacency matrix entries directly determine embedding distances. Taking expectation over the initial embeddings (assuming Gaussian initialization with zero mean), the conditional entropy can be approximated as:
$H\bigl{(}\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)}\bigr{)}\;\propto\;\sum_{j,k}\bigl{(}\hat{\mathbf{A}}^{l}_{up,j}-\hat{\mathbf{A}}^{l}_{p,k}\bigr{)}^{2}\mathbb{E}\bigl{[}\|\mathbf{x}_{j}^{(0)}-\mathbf{x}_{k}^{(0)}\|^{2}\bigr{]}.$ (14)
Crucially, as layer depth $l$ increases, the phenomenon of over-smoothing causes $\hat{\mathbf{A}}^{l}_{up,j}$ and $\hat{\mathbf{A}}^{l}_{p,k}$ to become increasingly similar for all node pairs (Eq. (9)). Consequently, as these adjacency matrix differences shrink, the squared embedding distances also decrease, directly causing a layer-wise monotonic reduction in conditional entropy:
$H\bigl{(}\mathbf{x}_{up}^{(l)}\mid\mathbf{x}_{p}^{(l)}\bigr{)}\;\geq\;H\bigl{(}\mathbf{x}_{up}^{(l+1)}\mid\mathbf{x}_{p}^{(l+1)}\bigr{)}.$ (15)
From this monotonic trend, we conclude that repeated propagation via higher-order adjacency matrices leads to embedding homogenization, which progressively reduces conditional entropy and thus weakens the effectiveness of supervised alignment at deeper GCN layers. Our theoretical findings not only explain why deeper GCN layers diminish alignment effectiveness but also highlight the importance of designing adaptive strategies that dynamically adjust alignment contributions across layers. This theoretical insight serves as a foundation for developing novel approaches to mitigate the impact of embedding homogenization in graph-based learning.
III-B Experimental Validation
(a) Conditional Entropy and Similarity
(b) Supervised Alignment
Figure 2: ((Left) Variation in conditional entropy and embedding similarity across GCN layers, illustrating how deeper layers lead to increased homogenization. (Right) Effect of applying supervised alignment at different GCN depths, highlighting its impact on recommendation performance and the diminishing gains at deeper layers.
We perform experiments on the Gowalla dataset to empirically examine the relationship among over-smoothing, embedding homogenization, and supervised alignment effectiveness in deeper layers. First, we randomly sample 1,000 pairs of popular and unpopular items and calculate their conditional entropy ($H(\mathbf{X}_{up}^{(l)}|\mathbf{X}_{p}^{(l)})$) and cosine similarity across different layers of LightGCN. As shown in Fig.2a, the conditional entropy between popular and unpopular items significantly decreases as the layer depth increases, coinciding with increased embedding similarity. This observation empirically confirms our theoretical finding that *embedding homogenization occurs due to the progressive uniformity of the adjacency matrix entries in deeper layers (Eq. (9)), which directly reduces the distinguishability of unpopular item embeddings.
Furthermore, we evaluate the effectiveness of supervised alignment using the PAAC method [19] applied at various layer depths on Gowalla, as illustrated in Fig.2b. Performance metrics (e.g., NDCG@20, Recall@20) clearly demonstrate that supervised alignment applied at shallow layers significantly improves recommendation quality, while alignment in deeper layers leads to negative gains. This aligns with our theoretical derivation that conditional entropy decreases monotonically with depth (Eq. (15)), leading to diminishing information transfer potential from popular to unpopular items. Specifically, as shown in Eq. (14), the reduced variance in adjacency matrix weights across nodes causes the expected embedding distance to shrink, directly driving the observed entropy decay.
In summary, both theoretical analysis and empirical validation underline the critical limitation of existing supervised alignment approaches: repeated propagation via higher-order adjacency matrices leads to embedding homogenization(Eq. (9)), which progressively reduces conditional entropy (Eq. (15)) and thus weakens the effectiveness of supervised alignment at deeper GCN layers. These findings strongly support the necessity of developing adaptive strategies that adjust alignment strength based on layer depth, as explored in our proposed dual adaptive strategy.
IV Methodology
Figure 3: The Graph-Structured Dual Adaptation Framework (GSDA)  consists of two key modules: (1) Adaptive Hierarchical Supervised Alignment, which reduces over-smoothing in deeper GCN layers; and (2) Adaptive Re-weighting Contrastive Learning, which dynamically adjusts sample weights based on real-time popularity distributions.
To address the limitations identified earlier, we propose a novel dual-adaptive framework, termed  Graph-Structured Dual Adaptation Framework (GSDA), which explicitly leverages the structural and distributional properties encoded within the graph’s high-order adjacency matrix to mitigate popularity bias. Our theoretical and empirical analyses uncover a critical yet previously overlooked limitation: as Graph Convolutional Networks (GCNs) deepen, embeddings derived from high-order adjacency matrices ($\hat{\mathbf{A}}^{l}$) progressively become more uniform due to the over-smoothing effect. This increasing uniformity results in significant homogenization of node embeddings, thereby reducing the conditional entropy between popular and unpopular item representations. Consequently, the effectiveness of supervised alignment strategies diminishes, particularly at deeper layers.
Inspired by these insights, we propose  Graph-Structured Dual Adaptation Framework (GSDA), which leverages structural features and distribution information embedded in the high-order adjacency matrix to address these challenges. The specific framework model diagram is shown in Fig.3. Specifically, we introduce two innovative modules: (1) Adaptive Hierarchical Supervised Alignment, which employs a carefully designed adaptive weight decay mechanism based on structural information from higher-order adjacency matrices to alleviate the weakening effects of supervised alignment in deeper layers; and (2) Adaptive Re-weighting Contrastive Learning, which dynamically estimates item popularity distributions utilizing adjacency matrix information, adaptively adjusting embedding weights in real-time. This eliminates the dependency on fixed hyperparameters, enhancing the model’s adaptability and robustness across varying item distributions.
IV-A Adaptive Hierarchical Supervised Alignment
Our theoretical and empirical analyses have clearly demonstrated that the effectiveness of supervised alignment between popular and unpopular items deteriorates significantly with increasing Graph Convolutional Network (GCN) layer depth. The fundamental reason behind this issue is the homogenization of node embeddings in deeper layers, caused by the over-smoothing effect when repeatedly propagating embeddings via the high-order normalized adjacency matrix ($\hat{\mathbf{A}}^{l}$). Specifically, as the layer depth $l$ increases, embeddings become increasingly uniform, reducing the conditional entropy between popular and unpopular item representations. This entropy reduction directly limits the ability of supervised alignment to transfer useful information from popular items to unpopular ones, ultimately weakening its effectiveness in mitigating popularity bias.
To address this issue, we propose the Adaptive Hierarchical Supervised Alignment module, which dynamically adjusts the alignment influence across different GCN layers based on structural characteristics encoded in the adjacency matrix. Following the embedding propagation mechanism used in LightGCN [23], node embeddings at layer $l$ are computed as:
$\mathbf{X}^{(l)}=\hat{\mathbf{A}}^{l}\mathbf{X}^{(0)},$ (16)
where $\mathbf{X}^{(0)}$ represents the initial node embeddings, and $\hat{\mathbf{A}}$ denotes the normalized adjacency matrix. At the base layer ($l=0$), the adjacency matrix is defined as the identity matrix, i.e., $\mathbf{A}^{0}=\mathbf{I}$, ensuring the embeddings at layer 0 solely reflect initial self-information:
$\mathbf{X}^{(0)}=\mathbf{I}\mathbf{X}^{(0)}=\mathbf{X}^{(0)}.$ (17)
Given our theoretical insight that deeper layers exhibit increased embedding uniformity, we introduce an adaptive layer-wise weighting strategy informed by the Frobenius norm of adjacency matrix powers:
$w_{l}=\frac{\|\hat{\mathbf{A}}^{l}\|_{F}}{\sum_{k=0}^{L}\|\hat{\mathbf{A}}^{k}\|_{F}},$ (18)
where $\|\hat{\mathbf{A}}^{l}\|_{F}$ represents the Frobenius norm of the adjacency matrix at layer $l$, quantifying the diversity or richness of structural information retained after $l$-hop propagation. The Frobenius norm effectively measures how distinctively the adjacency structure is preserved at each propagation step. As $l$ increases, the entries in $\hat{\mathbf{A}}^{l}$ tend to become more uniform due to over-smoothing, leading to a smaller Frobenius norm. Consequently, this adaptive weighting strategy inherently assigns greater emphasis to shallow layers (which preserve richer localized information) and progressively reduces weights at deeper layers.
Given the adaptive layer-wise weighting scheme, the hierarchical supervised alignment loss across layers is defined as:
$\mathcal{L}_{sa}=\frac{1}{L+1}\sum_{l=0}^{L}w_{l}\left(\frac{1}{|U|}\sum_{u\in\mathcal{U}}\frac{1}{|\boldsymbol{I}_{u}|}\sum_{i\in\boldsymbol{I}_{u}^{up},j\in\boldsymbol{I}_{u}^{p}}\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}\right),$ (19)
where $U$ denotes the set of users, and $\boldsymbol{I}_{u}^{p}$ and $\boldsymbol{I}_{u}^{up}$ are the sets of popular and unpopular items interacted with by user $u$, respectively. The term $\|\mathbf{x}_{i}^{(l)}-\mathbf{x}_{j}^{(l)}\|^{2}$ explicitly aligns embeddings of popular and unpopular items at each layer, with weights adaptively adjusted to reflect the structural significance captured at different depths.
By emphasizing shallow-layer embeddings, where conditional entropy and representational distinctiveness remain relatively high, our adaptive strategy preserves the beneficial effects of supervised alignment. At deeper layers, where embeddings suffer more pronounced homogenization, it systematically reduces alignment weights. This design allows  GSDA  to dynamically adjust alignment contributions based on structural insights derived directly from the graph. As a result, it enhances robustness against popularity bias and improves recommendation performance.
IV-B Adaptive Re-weighting Contrastive Learning
Existing contrastive learning methods in recommendation typically use fixed hyperparameters to balance the influence between popular and unpopular items, which restricts their adaptability and effectiveness across different datasets and dynamic environments. Furthermore, these static strategies neglect the nuanced variations in item popularity distributions, leading to potential misrepresentations and suboptimal performance. To explicitly overcome these limitations, we propose an adaptive re-weighting contrastive learning module that dynamically adjusts the weights of samples based on real-time item popularity distributions extracted directly from structural characteristics of the graph’s adjacency matrix.
Specifically, rather than assigning uniform or fixed weights, we introduce a user-specific weighting strategy that more accurately captures each user’s contribution to item popularity. Each user $u$ is assigned a weight inversely proportional to their degree (number of interactions), formally defined as:
$w_{u}=\frac{1}{d_{u}+\varepsilon},$ (20)
where $d_{u}$ is the degree of user $u$, and $\varepsilon$ is a small constant introduced to maintain numerical stability, preventing division by zero when users have minimal interactions. By introducing this user-specific weighting, the contributions from highly active users are balanced, ensuring a more accurate and nuanced estimation of item popularity.
Based on this adjusted weighting, we then derive the item popularity $p_{i}^{\ast}$ by aggregating interactions between users and items, weighted by the user-specific factors:
$p_{i}^{\ast}=\sum_{u\in\mathcal{U}}w_{u}\cdot\widetilde{A}_{u,i+M},$ (21)
where $\widetilde{A}_{u,i+U}$ represents the normalized interaction strength between user $u$ and item $i$. This formulation provides a refined and dynamic estimation of item popularity, effectively reflecting the true underlying distributional dynamics captured from the graph structure.
With this dynamically computed item popularity distribution, we quantify the level of imbalance between popular and unpopular items using the Gini coefficient, defined as:
$w=\frac{1}{2I\overline{p^{\ast}}}\sum_{i=0}^{I-1}\sum_{j=0}^{I-1}\left|p_{i}^{\ast}-p_{j}^{\ast}\right|,$ (22)
where $\overline{p^{\ast}}$ denotes the average weighted popularity of all items. The Gini coefficient measures the inequality or imbalance in the distribution of item popularity, with a higher value indicating a larger imbalance (i.e., few items dominate interactions).
Finally, we integrate this dynamic measurement into contrastive learning framework, adaptively adjusting the weights for popular and unpopular items based on the imbalance captured by the Gini coefficient (Eq. (22)). Specifically, we propose a variant of contrastive loss that dynamically modulates the influence of popular and unpopular items when used as positive or negative samples:
$\mathcal{L}_{cl}=(1-w)\mathcal{L}_{cl}^{p}+w\mathcal{L}_{cl}^{up},$ (23)
where $\mathcal{L}_{cl}^{p}$ and $\mathcal{L}_{cl}^{up}$ represent the contrastive losses computed separately for popular and unpopular items. By adaptively adjusting weights based on $w$, our method dynamically reduces the dominance of the over-represented group. When the distribution is highly skewed (large $w$), the model strategically reduces the influence of popular items to enhance embedding diversity and robustness. Conversely, when item popularity is relatively balanced (small $w$), the method treats both groups more evenly, maintaining optimal performance.
This adaptive design is driven entirely by real-time insights from the high-order adjacency matrix, eliminating reliance on fixed hyperparameters. By dynamically capturing nuanced graph-based structural signals, our approach ensures effective adaptation to varying item distributions, thereby significantly improving the robustness and generalizability of recommendations while effectively mitigating popularity bias.
IV-C Model Optimization and Algorithm
To effectively mitigate the two main limitations of popularity bias, we employ a multi-task training approach [18, 31]. The overall loss function for  GSDA  is defined as follows:
$\mathcal{L}=\mathcal{L}_{rec}\;+\;\lambda_{1}\,\mathcal{L}_{sa}\;+\;\lambda_{2}\,\mathcal{L}_{cl},$ (24)
where $\mathcal{L}_{rec}$ is the base recommendation loss, capturing fundamental user–item preference signals (e.g., BPR [14] or another ranking objective); $\mathcal{L}_{sa}$ is the adaptive hierarchical supervised alignment loss, designed to address the over-smoothing effect in deeper GCN layers and preserve discriminative item representations;$\mathcal{L}_{cl}$ is the adaptive re-weighting contrastive loss, derived from dynamic item popularity estimation, enabling the model to flexibly adjust sample weights based on real-time popularity distributions. The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ control the relative importance of $\mathcal{L}_{sa}$ and $\mathcal{L}_{cl}$, respectively. After training, we perform inference by computing the dot product between the final user and item embeddings, ranking items based on their predicted relevance [26]. The detailed algorithm for  GSDA  is presented in Algorithm 1.
Our theoretical and experimental findings have demonstrated that over-smoothing in deeper GCN layers—caused by high-order adjacency matrices becoming increasingly uniform—significantly undermines the effectiveness of supervised alignment. Meanwhile, fixed hyperparameter schemes in contrastive learning fail to accommodate diverse item popularity distributions, resulting in suboptimal performance. By integrating these insights into a graph-structured, dual-adaptive framework, GSDA systematically addresses both issues. First, by adapting layer-wise alignment weights to the structure of higher-order adjacency matrices, we reduce the detrimental impact of over-smoothing on supervised alignment. Second, by dynamically estimating item popularity from real-time adjacency matrix signals, we remove the reliance on static hyperparameters in contrastive learning. Through this synergy, GSDA not only boosts recommendation accuracy but also adapts robustly to varying popularity distributions across datasets. In doing so, it significantly alleviates the “rich-get-richer” phenomenon pervasive in recommender systems, offering a powerful and flexible solution to popularity bias.
Algorithm 1 Training Procedure of GSDA
0:  Interaction matrix $\mathcal{R}$, Normalized adjacency matrix $\widetilde{\mathbf{A}}$, Learning rate $\eta$, Hyperparameters $\lambda_{1},\lambda_{2}$, Layer depth $L$, Stability constant $\varepsilon$
0:  Trained model parameters $\Theta$
1:  Initialize $\Theta$ (user/item embeddings and GCN weights)
2:  for epoch $=1$ to $T$ do
3:     for mini-batch $\mathcal{B}\subseteq\mathcal{R}$ do
4:        Embedding Propagation:
5:        for $l=0$ to $L$ do
6:           $\mathbf{X}^{(l)}\leftarrow\widetilde{\mathbf{A}}^{l}\mathbf{X}^{(0)}$ (Eq. 16)
7:        end for
8:        Adaptive Hierarchical Alignment:
9:        for $l=0$ to $L$ do
10:           $w_{l}\leftarrow\|\widetilde{\mathbf{A}}^{l}\|_{F}/\sum_{k=0}^{L}\|\widetilde{\mathbf{A}}^{k}\|_{F}$ (Eq. 18)
11:        end for
12:        Compute $\mathcal{L}_{sa}$ (Eq. 19 )
13:        Adaptive Contrastive Reweighting:
14:        for each user $u\in\mathcal{B}$ do
15:           $w_{u}\leftarrow 1/(d_{u}+\varepsilon)$ (Eq. 20 )
16:        end for
17:        Compute $p_{i}^{*}\leftarrow\sum_{u}w_{u}\widetilde{A}_{u,i+U}$ (Eq. 21 )
18:        Compute $w\leftarrow\text{Gini}(\{p_{0}^{*},...,p_{I-1}^{*}\})$ (Eq. 22 )
19:        Compute $\mathcal{L}_{cl}\leftarrow(1-w)\mathcal{L}_{cl}^{p}+w\mathcal{L}_{cl}^{up}$ (Eq. 23 )
20:        Multi-task Optimization:
21:        $\mathcal{L}\leftarrow\mathcal{L}_{rec}+\lambda_{1}\mathcal{L}_{sa}+\lambda_{2}\mathcal{L}_{cl}$ (Eq. 24 )
22:        $\Theta\leftarrow\Theta-\eta\nabla_{\Theta}\mathcal{L}$Parameter update}
23:        if validation performance plateaus then
24:           break {Early stopping}
25:        end if
26:     end for
27:  end for
V Experiments
V-A Experimental Settings
V-A1 Datasets
We conducted experiments on three widely-used recommendation benchmarks: Gowalla [^1] http://snap.stanford.edu/data/loc-gowalla.html, Movielens-10M [^2] https://grouplens.org/datasets/movielens/, and Globo [^3] https://paperswithcode.com/dataset/news-interactions-on-globo-com. Gowalla is a location-based social networking dataset that captures sparse user-item interactions with a pronounced long-tail item distribution, ideal for evaluating popularity bias mitigation. Movielens-10M is a popular dataset comprising 10 million user-movie ratings, which we converted from explicit feedback to implicit feedback for consistency in our experiments. Globo is a news recommendation dataset capturing user interactions with rapidly evolving news articles, testing a model’s ability to adapt to dynamic item distributions. To ensure data quality, we followed established preprocessing practices [17, 10, 32], retaining only users and items with at least 10 interactions. Detailed dataset statistics are summarized in Tab.I.
TABLE I: The statistics of three datasets.
Datasets #Users #Itmes #Interactions Density
Gowalla 29,858 40,981 1,027,370 0.084%
ML-10M 69,166 8,790 5,000,415 0.823%
Globo 158,323 12,005 2,484,192 0.131%
V-A2 Baselines and Evaluation
To evaluate the effectiveness of  GSDA  in mitigating popularity bias, we implemented it using the state-of-the-art LightGCN framework [23]. We compared  GSDA  against several established debiasing baselines, including re-weighting models such as $\gamma$-AdjNorm [29] and DR-GNN [33], decorrelation approaches like MACR [10] and InvCF [34], as well as contrastive learning-based methods, including SGL [18], Adap-$\tau$ [35], and SimGCL [17]. Additionally, we benchmarked our method against the supervised alignment approach PAAC [19].
It is worth noting that while other works have explored mitigating popularity bias by regularizing graph structural information [28, 36], these approaches were excluded due to the unavailability of their implementation. Instead, we discuss their contributions in the related work section. This selection of baselines provides a comprehensive comparison across different debiasing paradigms, demonstrating the effectiveness of  GSDA  in addressing popularity bias.
To accurately evaluate the model’s ability to mitigate popularity bias, we adopted an unbiased dataset with uniformly distributed items in the test set, following established practices [37, 34, 28, 10]. This ensures that the evaluation focuses on the model’s effectiveness rather than its reliance on item popularity. Specifically, each item in the test set was assigned a fixed number of interactions, representing approximately 10% of the total dataset. To maintain the integrity of the test distribution, an additional 10% of interactions were randomly allocated to the validation set, while the remaining interactions were used for training. We evaluated Top-$K$ recommendation performance using $Recall@K$, $HR@K$, and $NDCG@K$. $Recall@K$ and $HR@K$ measure the model’s ability to provide comprehensive coverage, while $NDCG@K$ focuses on the quality of item ranking. To ensure fairness and avoid selection bias, we employed a full ranking strategy [38, 39], which evaluates all items rather than a subset [26]. Each experiment was repeated five times with different random seeds to account for variability, and the average scores were reported [26, 13]. This robust evaluation protocol ensures reliable and unbiased assessment of the model’s performance in mitigating popularity bias.
V-A3 Hyper-Parameter Settings.
We initialize the model parameters using the Xavier initializer [40] and optimize the model with the Adam optimizer [41] at a learning rate of 0.001. The embedding size is set to 64. For all datasets, the batch size is fixed at 2048, and the $L_{2}$ regularization coefficient $\lambda_{3}$ is set to 0.0001. Motivated by theoretical insights regarding sensitivity to alignment and contrastive losses, we systematically searched hyperparameters $\lambda_{1}$ (1 to 1000) and $\lambda_{2}$ (0.1 to 20) for optimal adaptability across datasets.
V-B Overall Performance
TABLE II: Performance comparison across three public datasets with $K=20$. The best results are highlighted in bold, and the second-best are underlined. Superscripts * denote $p\leq 0.05$ from a paired t-test comparing  GSDA  to the top-performing baseline (relative improvements are indicated as Imp.).
Model Gowalla ML-10M Globo
$Recall@20$ $HR@20$ $NDCG@20$ $Recall@20$ $HR@20$ $NDCG@20$ $Recall@20$ $HR@20$ $NDCG@20$
MF 0.0343 0.0422 0.0280 0.1283 0.1424 0.1026 0.0803 0.0806 0.0418
LightGCN 0.0380 0.0468 0.0302 0.1650 0.1791 0.1255 0.1233 0.1329 0.0745
$\gamma$-Adjnorm 0.0328 0.0409 0.0267 0.1103 0.1235 0.0881 0.1166 0.1169 0.0573
DR-GNN 0.0972 0.1048 0.0639 0.1490 0.1656 0.1318 0.1460 0.1465 0.0817
MACR 0.0908 0.1086 0.0600 0.1302 0.1591 0.1211 0.1275 0.1302 0.0635
InvCF 0.1001 0.1202 0.0662 0.1593 0.1725 0.1283 0.1469 0.1492 0.0804
SGL 0.0.114 0.1213 0.0778 0.1741 0.1893 0.1356 0.1522 0.1526 0.0829
Adap-$\tau$ 0.1182 0.1248 0.0794 0.1705 0.1802 0.1299 0.1562 0.1574 0.0836
SimGCL 0.1194 0.1228 0.0804 0.1720 0.1875 0.1348 0.1585 0.1589 0.0849
PAAC 0.1232 0.1321 0.0848 0.1791 0.1959 0.1458 0.1610 0.1615 0.0862
GSDA 0.1300* 0.1383* 0.0882* 0.1886* 0.2086* 0.1515* 0.1668* 0.1675* 0.0908*
Imp. +5.52% +4.69% +4.01% +5.30% +6.48% +3.91% +3.60% +3.71% +5.33%
As shown in Tab.II, we compare  GSDA  with several debiasing baselines across three datasets. The best-performing results are highlighted in bold, and the second-best results are underlined.  GSDA  consistently outperforms all baselines across every metric and dataset, achieving significant improvements. From the experimental results, we observe the following key insights:
  • GSDA achieves substantial improvements compared to competitive baselines such as PAAC, Adap-$\tau$, and SimGCL, with performance gains ranging from 3.53% to 6.99%. For example, on the Movielens-10M dataset,  GSDA outperforms the best-performing baseline PAAC by 3.91% in NDCG@20 and by 5.30% in Recall@20. Similarly, on the Globo dataset,  GSDA achieves a 3.71% improvement in HR@20. Although PAAC demonstrates strong performance due to its supervised alignment strategy, which transfers information from popular items to unpopular ones,  GSDA surpasses PAAC by incorporating dual adaptive mechanisms. Specifically, our adaptive hierarchical alignment better captures distinct features at different layers, while our adaptive re-weighting contrastive learning dynamically adjusts sample weights, leading to superior representation learning.
  • Contrastive learning-based baselines such as SGL and SimGCL show reasonable performance by promoting embedding uniformity and mitigating popularity bias. In contrast, decorrelation-based methods like MACR and InvCF, which explicitly attempt to separate item embeddings from popularity signals, exhibit inconsistent results. This indicates that popularity is not inherently harmful and may even signal item quality. Notably, PAAC achieves strong results because its supervised alignment effectively leverages information transfer from popular to unpopular items, enhancing their embeddings.
  • Our model  GSDA  is particularly effective in addressing data sparsity. For instance, on the Gowalla dataset, characterized by sparse user-item interactions,  GSDA achieves a notable 4.01% improvement in NDCG@20 and a 5.52% increase in Recall@20 over the strongest baseline. These improvements are attributed to our adaptive hierarchical alignment, which prioritizes shallow layers that retain distinct embedding features, and our adaptive re-weighting module, which dynamically balances contributions across items with varying levels of popularity. Collectively, these strategies ensure robust and discriminative embeddings, even in sparse data environments.
  • The dual adaptive mechanisms in  GSDA are essential for its superior performance. The adaptive hierarchical supervised alignment addresses the over-smoothing issue commonly observed in deeper GCN layers by adaptively modulating the alignment contributions from each layer, thereby preventing embedding homogenization. Meanwhile, the adaptive re-weighting contrastive learning component dynamically adjusts the contrastive weights based on batch-level item popularity distributions, informed by user-specific interactions. This dual adaptation strategy not only effectively mitigates popularity bias but also enhances model generalizability and robustness across datasets with diverse interaction patterns and varying levels of sparsity, eliminating reliance on extensive hyperparameter tuning.
V-C Ablation Study
TABLE III: An ablation study of  GSDA  on the Gowalla dataset, with the best-performing models for each metric highlighted in bold.
Model $Recall@20$ $HR@20$ $NDCG@20$
GSDA 0.1311 0.1391 0.0884
GSDA-w/o SA 0.1240 0.1292 0.0823
GSDA-w/o CL 0.1206 0.1287 0.0820
GSDA-SAF 0.1252 0.1348 0.0858
GSDA-SA0 0.1264 0.1353 0.0872
GSDA-FRW 0.1288 0.1377 0.0874
To clearly demonstrate the effectiveness of each model component, we conducted ablation experiments on the Gowalla dataset, which is notably sparser and more representative of real-world recommendation scenarios. Similar patterns and conclusions were observed on other datasets. Tab.III presents the Top-20 recommendation performance of  GSDA and several carefully designed variants. Specifically, we introduce the following four variants:
  • GSDA-w/o SA: This variant entirely removes the adaptive hierarchical supervised alignment loss, thus eliminating the explicit alignment between popular and unpopular items. This setting isolates the contribution of the adaptive contrastive module, clarifying its independent effect on popularity bias mitigation.
  • GSDA-w/o CL: This variant omits the adaptive re-weighting contrastive loss, focusing solely on adaptive hierarchical supervised alignment. This setup allows us to independently assess how the hierarchical alignment contributes to enhancing item embedding distinctiveness.
  • GSDA-SAF: This variant replaces the adaptive layer-specific alignment strategy with a single fused alignment. Specifically, item embeddings aggregated from all GCN layers are aligned as a single unified representation rather than aligning embeddings layer-by-layer. This evaluates whether explicit layer-wise distinctions provide advantages over fused representations.
  • GSDA-SA0: This variant limits the supervised alignment process exclusively to the initial embedding layer ($\mathbf{h}_{i}^{0}$). By restricting alignment to the shallowest layer, we test the significance of incorporating richer structural information available in deeper layers.
  • GSDA-FRW: This variant employs fixed contrastive learning weights instead of dynamically computing them via the Gini coefficient. Static weights disregard the adaptive distribution of item popularity, allowing us to examine the necessity and impact of our adaptive weighting scheme.
From the ablation results summarized in Tab.III, we draw several insightful conclusions:
  • The complete  GSDA model consistently achieves superior results across all metrics, outperforming the best variant (GSDA-FRW) by 1.79%, 0.98%, and 1.37% in Recall@20, HR@20, and NDCG@20, respectively. This clearly demonstrates the complementary nature and effectiveness of our dual-adaptive design leveraging both graph structure and distributional characteristics.
  • Excluding the adaptive hierarchical supervised alignment (GSDA-w/o $SA$) results in a significant drop in performance (5.42% reduction in Recall@20 compared to the complete model). This highlights the critical role of hierarchical alignment in preserving embedding distinctiveness across layers, particularly combating the homogenization caused by over-smoothing.
  • Variants utilizing fused representations (GSDA-SAF) or aligning solely the initial layer (GSDA-SA0) exhibit inferior performance compared to the full hierarchical model, underscoring that adaptive layer-wise alignment better captures structural nuances and enhances embeddings more effectively than a simplified or shallow alignment strategy.
  • Removing adaptive contrastive weighting (GSDA-w/o $CL$) or using fixed contrastive weights (GSDA-FRW) clearly reduces recommendation quality, emphasizing that dynamic adjustment based on the Gini coefficient effectively manages popularity distributions, preventing overemphasis on popular items and ensuring balanced representation.
Collectively, these experiments validate that our dual-adaptive framework, leveraging both the hierarchical structural information of the graph and dynamic popularity distributions, is essential for effectively mitigating popularity bias, improving recommendation performance, and ensuring robust generalization across diverse dataset scenarios.
V-D Hyper-Parameter Sensitivities
In this section, we performed a comprehensive analysis of hyper-parameter sensitivities to validate and further understand the effectiveness of our proposed model. Specifically, we examined the influence of the alignment and contrastive learning coefficients ($\lambda_{1}$ and $\lambda_{2}$), the adaptive weighting parameter ($w$), and the impact of different grouping ratios for item popularity. Additionally, we analyzed the relationship between the number of hyperparameters and model performance, highlighting that our proposed adaptive strategies achieve superior recommendation accuracy with fewer hyperparameters compared to existing methods. Throughout our experiments, the dashed lines in each figure consistently denote the performance achieved by the strongest baseline. The results emphasize the importance of carefully balancing the adaptive hierarchical supervised alignment and adaptive re-weighting contrastive learning components, as well as the clear advantage of dynamically adjusting weights according to real-time item popularity distributions. Note that we primarily present experimental results on the Gowalla dataset due to its sparse interactions and representative nature. Our empirical findings remain consistent across other datasets.
V-D1 Effect of $\lambda_{1}$ and $\lambda_{2}$
As shown in Eq. (24), the hyperparameters $\lambda_{1}$ and $\lambda_{2}$ control the relative contributions of the adaptive hierarchical supervised alignment loss and the adaptive re-weighting contrastive loss, respectively. Fig.4 illustrates the impact of these hyperparameters on the model’s performance for the Gowalla dataset, with the dashed line indicating the performance level of the best-performing baseline.
In the left subfigure, performance initially improves as $\lambda_{1}$ increases, underscoring the effectiveness of adaptive hierarchical supervised alignment in enhancing the embeddings of unpopular items. However, performance gains gradually diminish after surpassing an optimal point and subsequently decline. This decline is attributed to an excessive emphasis on unpopular items, which introduces noise and reduces the overall recommendation quality.
Similarly, in the right subfigure, moderately increasing $\lambda_{2}$ yields performance improvements by promoting consistency among item embeddings and effectively mitigating popularity bias. Nevertheless, excessively large values of $\lambda_{2}$ result in either stagnation or deterioration in performance. This phenomenon occurs because overly strict regularization on representation uniformity can impair the model’s capacity to learn personalized item preferences. These observations collectively highlight the necessity of balancing the supervised alignment and contrastive components to achieve optimal recommendation performance while effectively addressing popularity bias.
Figure 4: Comparison of recommendation performance ($NDCG@20$ and $HR@20$) for varying hyperparameters $\lambda_{1}$ and $\lambda_{2}$ against the optimal baseline on the Gowalla dataset, illustrating the sensitivity and effectiveness of our adaptive modules.
Figure 5: (Left) Comparison of dynamic and static weighting parameters in adaptive contrastive learning on the Gowalla dataset, highlighting the effectiveness of dynamic adaptation. (Right) Recommendation performance under varying item grouping ratios on the Gowalla dataset, demonstrating the robustness of our adaptive strategy across different popularity distributions.
V-D2 Effect of Adaptive Weighting ($w$)
To further validate the effectiveness of our adaptive weighting strategy ($w$), we conducted comparative experiments on the Gowalla dataset, evaluating our dynamically computed weighting scheme (denoted as ours) against several fixed weighting values ranging from 0.1 to 1. The red line shows the recommendation performance achieved by  GSDA. The results presented in Fig.5a clearly demonstrate that our adaptive weighting approach consistently outperforms all fixed-weight counterparts, especially in terms of $Recall@20$ and $NDCG@20$. These improvements underscore the importance of dynamically adjusting contrastive weights based on real-time item popularity distributions, rather than relying on static hyperparameters.
Specifically, our adaptive weighting mechanism utilizes the Gini coefficient derived from the normalized adjacency matrix, integrating user-specific weighting to capture heterogeneous interaction patterns across users. This adaptive design reduces the need for extensive hyperparameter tuning, thus simplifying model optimization. Moreover, by dynamically responding to varying levels of data sparsity and shifting popularity distributions, the method significantly enhances the model’s robustness and adaptability. These findings confirm that dynamically adjusting the contrastive weighting parameter is essential for effectively mitigating popularity bias and achieving consistently superior recommendation performance.
V-D3 Effect of Group Ratio
To investigate the influence of the partition ratio between popular and unpopular items on recommendation performance, we experimented with dynamically categorizing the items within each mini-batch based on their relative popularity. Instead of using a fixed global popularity threshold, we classified the top $x\%$ items in terms of popularity as popular and the remaining $(100-x)\%$ as unpopular. While a 50% ratio was initially selected as a default setting following previous studies [19], we further explored varying this ratio from 10% to 90% in increments of 10%.
The results shown in Fig.5b demonstrate that mid-range ratios (between 40% and 60%) consistently yield superior recommendation performance, with the optimal results achieved at the 50% partition. Conversely, extreme ratios (e.g., 10%-30% or 70%-90%) led to significantly poorer performance, as these ratios either excessively focused the model’s attention on a limited set of popular items or overly emphasized unpopular items, hindering effective learning. These findings underscore the importance of maintaining a balanced representation of popular and unpopular items, highlighting the potential advantages of employing adaptive partitioning strategies tailored to the underlying popularity distribution for improving recommendation accuracy.
V-D4 Effect of Hyperparameter Complexity
Figure 6: Comparison of model complexity and recommendation effectiveness. The horizontal axis denotes the number of hyperparameters required by each model, while the vertical axis represents their recommendation performance measured by NDCG@20 on the Gowalla dataset.
To further highlight the practical advantages of our dual-adaptive strategy, we compared  GSDA to the baselines in terms of both performance (NDCG@20) and the number of required hyperparameters, as depicted in Figure 6. Although PAAC achieves strong performance, it relies on a higher number of hyperparameters, increasing the complexity of model tuning and limiting practical usability. In contrast, our proposed  GSDA, benefiting from its adaptive hierarchical alignment and adaptive re-weighting mechanisms, achieves superior recommendation performance with significantly fewer hyperparameters. This not only simplifies hyperparameter tuning but also enhances the robustness and generalizability of our model across diverse datasets and scenarios, underscoring the practical benefits of adaptive parameterization strategies.
V-E Debias Ability.
Figure 7: Comparative Analysis of Recommendation Performance Across Different Item Popularity Groups. This figure highlights the variations in performance metrics, specifically $NDCG@20$ and $HR@20$, across different item popularity groups on the Gowalla dataset, showcasing the impact of popularity bias on recommendation effectiveness.
To comprehensively evaluate our model’s debiasing capability, we analyze recommendation performance across five distinct item popularity groups (G1-G5, ordered by ascending popularity) on the Gowalla dataset, as illustrated in Fig. 7. These groups are partitioned based on interaction frequency quintiles, where G1 represents the least popular 20% of items and G5 the most popular 20%. We compare our model against baseline methods (LightGCN, SimGCL, PAAC) using NDCG@20 and HR@20. In Fig. 7, the line plot represents the number of items within each group, providing insight into the distribution of item popularity.
  • GSDA achieved notable performance gains for the less popular item groups (G1-G2). Specifically,  GSDA improved NDCG@20 by 12.7% for G1 and 9.3% for G2 compared to the best baseline (SimGCL). These improvements demonstrate the efficacy of our adaptive hierarchical alignment mechanism, which effectively transfers semantic knowledge from popular to unpopular items, enhancing the discriminability of long-tail item representations while preserving their intrinsic interaction characteristics.
  • Our approach significantly narrowed the performance gap between the least and most popular items. As illustrated in Fig. 7, the performance gradient from G1 to G5 was notably flattened, indicating a reduction in popularity-induced representation bias. This can be attributed to our adaptive re-weighting contrastive learning, which explicitly mitigates feature divergence across different popularity groups while retaining group-specific information.
  • Although the gains in the most popular item group (G5) were comparatively modest (+1.2% NDCG@20 over SimGCL),  GSDA achieved the highest overall NDCG@20 (+8.4%) across all groups. Crucially, the majority of the performance improvements (78%) came from enhancements in the long-tail groups (G1-G3), demonstrating that improving recommendation fairness for unpopular items simultaneously benefits overall recommendation quality.
VI Related Work
GCN-based Recommender Systems. Graph Convolutional Networks (GCNs) have significantly advanced Collaborative Filtering (CF) by capturing high-order user-item interactions [24, 42]. Unlike traditional CF models, which primarily focus on first-order user-item relationships, GCN-based methods propagate and aggregate information through graph structures, enabling more expressive user and item representations.
Various GCN-based recommendation models have been proposed to improve effectiveness and efficiency. NGCF [24] explicitly models high-order connectivity by stacking multiple graph convolutional layers, demonstrating the importance of propagating user-item signals beyond direct interactions. PinSage [43] adopts an efficient random-walk-based neighbor selection to scale GCN-based recommendations to large-scale graphs. LightGCN [23] simplifies GCN architectures by removing feature transformation and non-linear activations, focusing solely on neighbor aggregation to enhance both performance and efficiency. LR-GCCF [25] introduces a residual learning mechanism to mitigate the over-smoothing problem in deep GCN layers.
Despite these advancements, GCN-based methods still face challenges such as over-smoothing, where node representations become too similar as layers deepen, reducing model effectiveness [28, 29]. To address this, recent work explores residual connections [25], adaptive layer aggregation [23], and contrastive learning techniques [18] to enhance representation distinctiveness. However, balancing high-order propagation while maintaining feature diversity remains an open problem.
Supervised Alignment for Mitigating Popularity Bias. Supervised alignment in CF focuses on learning user and item representations based on observed interactions [13, 15, 44, 45, 14, 42]. Traditional alignment methods, such as BPR [14], optimize user-item ranking pairs to improve recommendation accuracy. However, these methods often struggle with biased representations, particularly for unpopular items with sparse interactions [19].
Recent approaches address this by leveraging group-level alignment. Popularity-aware alignment [19] aligns items interacted with by the same users, allowing unpopular items to inherit characteristics from popular ones. Other techniques, such as DebiasCF [46], integrate supervised signals with contrastive learning to enhance representation robustness against bias. While these methods improve alignment, they often apply uniform constraints across all GCN layers, failing to consider the varying effects of deeper layers. Our approach introduces an adaptive alignment strategy that dynamically adjusts alignment strength per layer, improving representation learning while mitigating popularity bias [47, 48, 42].
Re-weighting for Mitigating Popularity Bias. Re-weighting strategies address popularity bias through static weighting, reinforcement learning, and contrastive learning. (1) Static Weighting Strategies assign fixed weights to items based on their popularity, aiming to counteract the tendency of recommendation models to favor popular items [49, 11, 50]. Methods like inverse propensity scoring (IPS) [51, 52] estimate item exposure probabilities and re-weight interactions accordingly, while MACR [10] employs counterfactual inference to down-weight frequently interacted items [53]. However, these approaches require careful calibration and struggle with dynamic shifts in popularity distributions. (2) Reinforcement Learning-based Approaches dynamically adjust item weights by treating the recommendation process as a decision-making problem [54, 32]. Neural Selection (NSE) [55] employs an actor-critic framework to optimize item selection, while DRSA [56] models weight assignment as a Markov Decision Process. These approaches improve adaptability but introduce high computational costs and instability in sparse data scenarios. (3) Contrastive Learning Variants aim to balance representation learning by treating popularity bias as a distributional issue [57, 58, 59, 60, 18]. Methods like Self-supervised Graph Learning (SGL) [18] augment training by randomly removing edges, while SimGCL [17] injects controlled noise to promote uniformity. However, such strategies may inadvertently amplify bias by enforcing uniform sampling across items.
VII Conclusion
In this paper, we first provided both theoretical and empirical evidence demonstrating that repeated propagation through higher-order adjacency matrices in Graph Convolutional Networks (GCNs) leads to embedding homogenization. This phenomenon progressively reduces conditional entropy between popular and unpopular items, thereby weakening the effectiveness of supervised alignment strategies at deeper layers. Motivated by these insights, we introduced  GSDA, a novel dual-adaptive approach explicitly designed to mitigate popularity bias by effectively exploiting graph structural information and dynamically adapting to item popularity distributions. Specifically, we proposed an adaptive hierarchical supervised alignment module, which dynamically adjusts the contribution from each GCN layer to counteract the over-smoothing problem inherent in deeper layers. Complementing this, we introduced an adaptive re-weighting contrastive learning module, which dynamically adjusts the weights of item representations based on their real-time popularity distributions. This dual-adaptive design enables  GSDA to flexibly respond to varying item popularity patterns without depending on static hyperparameters.